OpenStudy (anonymous):

how do i find the equations of the tangent lines to the curve 5X^2 - y^2 = 41 at the point (3,-2)..its urgent cuz i got a test in like 4 hours..

OpenStudy (anonymous):

@andesa , u know how to differenciate?

OpenStudy (anonymous):

let \[y_0 = -2 \text{ , } x_0 = 3\] the equation of your line will be: \[y - y_0 = m(x - x_0)\] differentiating your curve eq. to find m: \[5x^2 - y^2 = 41\] \[10x - 2y\frac{dy}{dx} = 0\] \[\frac{dy}{dx} = \frac{10x}{2y} = \frac{5x}{y}\] at your point we have \[m = \frac{dy}{dx} = \frac{5(3)}{-2} = \frac{-15}{2}\]

OpenStudy (anonymous):

this requires a bit of knowledge of implicit differentiation, are you ok with that?

OpenStudy (anonymous):

yes!..thanks a lot!!

OpenStudy (anonymous):

well, after differenciating, u will get: 10x-2y dy/dx=0 dy/dx= 10x/2y= 5x/y slope of the tangent at any point (x,y)= 5x/y u have to put the value of the given point, so, slope=5*3/(-2)=-15/2 now, equation of the tangent: y=-15/2 x+c as the line passes through (3,-2).. -2=-15/2 *3+c u will get the value of c, u will get the answer